14. Say whether each of the following is about 1%, 10%, 25%, or 50%- (b) 99 out of 407 (d) 99 out of 197 (a) 39 out of 398 (c) 57 out of 209
Answer #1

Answer

(A) 39 out of 398 means 39/398 = 0.098

convering it into %, we get 0.098*100 = 9.8%, i.e. equal to 10%

So, 39 out of 398 is about 10%

(B) 99 out of 407 means 99/407 = 0.2432

convering it into %, we get 0.2432*100 = 24.32%, i.e. equal to 25%

So, 99 out of 407 is about 25%

(C) 57 out of 209 means 57/209 = 0.2727

convering it into %, we get 0.2727*100 = 27.27%, i.e. equal to 25%

So, 57 out of 209 is about 25%

(D) 99 out of 197 means 99/197 = 0.5025

convering it into %, we get 0.5025*100 = 50.25%, i.e. equal to 50%

So, 99 out of 197 is about 50%